Calculate the final specific heat capacity (SHC) of the mixture of 30g Nacl (c=867 j/kgk) and of 30g Nacl(c=867j/kgk) and 5g of kcl (c=682 j/kgk).

1
Expert's answer
2020-11-30T14:52:14-0500

The final specific heat capacity (SHC) of the mixture:


"C=m_1c_1+m_2c_2\\C=(0.03)(867)+(0.005)(682)=29.4\frac{J}{K}"